# RefDrone: A Challenging Benchmark for Referring Expression Comprehension in Drone Scenes

### v1

@@ Abstract
Drones have become prevalent robotic platforms with diverse applications, showing significant potential in Embodied Artificial Intelligence (Embodied AI). Referring Expression Comprehension (REC) enables drones to locate objects based on natural language expressions, a crucial capability for Embodied AI. Despite advances in REC for ground-level scenes, aerial views introduce unique challenges including varying viewpoints, occlusions and scale variations. To address this gap, we introduce RefDrone, a REC benchmark for drone scenes. RefDrone reveals three key challenges in REC: 1) multi-scale and small-scale target detection; 2) multi-target and no-target samples; 3) complex environment with rich contextual expressions. To efficiently construct this dataset, we develop RDAgent (referring drone annotation framework with multi-agent system), a semi-automated annotation tool for REC tasks. RDAgent ensures high-quality contextual expressions and reduces annotation cost. Furthermore, we propose Number GroundingDINO (NGDINO), a novel method designed to handle multi-target and no-target cases. NGDINO explicitly learns and utilizes the number of objects referred to in the expression. Comprehensive experiments with state-of-the-art REC methods demonstrate that NGDINO achieves superior performance on both the proposed RefDrone and the existing gRefCOCO datasets. The dataset and code will be publicly at https://github.com/sunzc-sunny/refdrone.
@@ 1 Introduction
Table 1: Comparison of referring expression comprehension datasets: RefCOCO/+/g [45, 29], gRefCOCO [22], D3 [42], RIS-CQ [13], RSVG [47], and RefDrone (ours). Avg. object: average number of objects per expression. Avg. length: words per expression. No target: expressions without referred objects. Expression type: expression generation method. Small target: small-scale target detection challenge.
RefCOCO/+/g [45, 29] gRefCOCO [22] D3 [42] RIS-CQ [13] RSVG [47] RefDrone Image source COCO [20] COCO [20] COCO [20] VG [16]+COCO [20] DIOR [18] VisDrone [49] Avg. object 1.0 1.4 1.3 3.6 2.2 3.8 Avg. length 3.6/3.5/8.4 4.9 6.3 13.2 7.5 9.0 No target Expression type Manual Manual Manual LLM (GPT3.5-t) Templated LMM (GPT4-o) Small target
@@
Drones/UAVs have become increasingly popular in our daily lives, serving both personal and professional purposes, such as entertainment, package delivery, traffic surveillance, and emergency rescue. Their ability to move freely and their broad observation capabilities make them important platforms for Embodied AI [6, 27, 17, 8, 25]. A crucial capability in Embodied AI is Referring Expression Comprehension (REC) [35, 40, 2, 39], which serves as a critical bridge between natural language understanding and visual perception. REC requires drones to localize specific objects in images based on natural language expressions. However, existing REC datasets primarily focus on ground-level perspectives, such as the RefCOCO [45, 29] dataset. The application of REC in drone-based scenarios presents unique challenges, including extreme viewpoint variations, occlusions, and scale variations across objects.
@@
In this work, we introduce RefDrone, a challenging benchmark designed for REC in drone scenes. The RefDrone dataset contains 17,900 referring expressions across 8,536 images, comprising 63,679 object instances. As illustrated in Figure 1, RefDrone poses three key challenges that distinguish it from existing datasets: (1) multi-scale and small-scale target detection, with 31% small objects and 14% large objects; (2) multi-target and no-target samples, where expressions can refer to any number of objects from 0 to 242; and (3) complex environment with rich contextual expressions. The environmental complexity manifests in varied viewpoints, diverse lighting conditions, and complex backgrounds. The contextual richness is reflected in expressions that describe spatial relations, attributes, and interactive relationships.
To efficiently construct this comprehensive dataset, we develop RDAgent (referring drone annotation framework with multi-agent system), a semi-automated annotation pipeline. RDAgent restructures traditional annotation workflows into an interactive system of multiple agents and human annotators. In this framework, agents with diverse roles collaborate through feedback loops to generate and verify annotations, while human involvement is minimized to quality control and minor adjustments. RDAgent significantly reduces annotation costs while maintaining high-quality standards for complex expressions, and it can be extended to generate annotations for other REC tasks.
@@
Furthermore, we propose a novel method called Number GroundingDINO (NGDINO). Our key insight is that utilizing the number information of referred objects enhances the handling of multi-target and no-target samples. NGDINO comprises three key components: (1) a number prediction head that estimates target object quantities, (2) a set of learnable number-queries encoding numerical patterns across different quantity samples, and (3) a number cross-attention module that integrates number queries with detection queries. Extensive experiments on both our drone-specific RefDrone benchmark and the general-domain gRefCOCO [22] dataset reveal that leveraging the number information greatly improves performance on multi-target and no-target samples.
@@
  1. 1.
    RefDrone Benchmark: The first comprehensive benchmark for referring expression comprehension in drone scenes. RefDrone poses three key challenges and provides comprehensive baseline evaluations across both specialized models and large multimodal models.
  2. 2.
    RDAgent Annotation Framework: A novel semi-automated annotation framework employs a multi-agent system. RDAgent reduces annotation costs while ensuring high-quality referring expressions.
  3. 3.
    NGDINO Method: A novel method specifically designed to handle multi-target and no-target samples through three components: a number prediction head, learnable number-queries, and number cross-attention. NGDINO achieves superior performance on both RefDrone and gRefCOCO datasets.
@@ 2.1 Referring expression understanding datasets
Referring expression understanding identifies specific regions within images or videos using natural language expressions. Two primary subtasks are Referring Expression Comprehension (REC) and Referring Expression Segmentation (RES). REC outputs detection bounding boxes, while RES outputs segmentation masks. Early datasets such as ReferIt [15] and RefCOCO [45] are pioneering but are limited to single-target expressions and simple expressions. Subsequently, datasets [22, 37, 13, 42] introduce more challenges. gRefCOCO [22] supports multi-target expressions. OV-VG [37] enables an open-vocabulary setting. D3 [42] and RIS-CQ [13] introduce more complex expressions. Additionally, domain-specific datasets [47, 30, 32] emerge to address unique applications. RSVG [47] focuses on remote sensing, RAVAR [30] targets video action recognition, and RIO [32] addresses affordance detection. However, drone scenes remain unexplored in REC tasks, which motivates our introduction of the RefDrone benchmark.
Recent advances in dataset creation have been significantly influenced by large language models (LLMs). Notable works such as LLaVA [24] and Ferret [44] show the effectiveness of LLMs in generating instruction tuning data. Similarly, RIS-CQ [13] and RIO [32] employ LLMs to generate complex referring expressions. However, these approaches have two limitations. First, LLMs are used mainly as text generators, potentially producing expressions without proper grounding in visual context. Second, the lack of iterative feedback mechanisms in the annotation pipelines may degrade annotation quality.
Figure 2: The overview of the RefDrone annotation process with RDAgent. Multiple specialized agents collaborate both with each other and human annotators through iterative feedback loops to generate high-quality annotations.
@@ 2.2 Referring expression comprehension methods
Referring expression comprehension methods can be broadly categorized into specialist models and large multimodal models (LMMs). Specialist models include two-stage and one-stage methods. Two-stage methods [12, 23, 10, 36, 9] typically approach REC as a ranking task, first generating region proposals and then ranking them based on language input. However, they are often criticized for slow inference speeds. In contrast, one-stage methods [14, 19, 26, 7, 43, 28] directly predict target regions guided by language. These approaches leverage transformer architectures to enable cross-modal interactions between visual and textual features. Among these methods, GroundingDINO (GDINO) [28] has gained widespread attention for its impressive results in both open-set object detection and REC tasks. However, GDINO lacks explicit mechanisms to handle multi-target or no-target scenarios.
Recent studies have increasingly evaluated LMMs [24, 44, 3, 5, 1, 31, 38, 21, 48, 41] on REC tasks to assess their visual-language understanding capabilities. These models leverage large-scale referring instruction tuning data to achieve remarkable performance. Despite their broad capabilities, LMMs consistently struggle with small object detection, primarily due to their inherent input resolution constraints. This constraint forces image downsampling, resulting in the loss of fine-grained visual details.
@@ 3.1 Data source
The RefDrone dataset builds upon VisDrone2019-DET [49], a drone-captured dataset for object detection. The source images are collected across multiple scenarios, illumination conditions, and flying altitudes. To ensure meaningful visual content, we implement two filtering criteria. At the image level, we exclude images containing less than 3 objects. At the object level, we exclude objects with bounding box areas below 64 pixels. The original VisDrone2019-DET annotations provide object categories and bounding box coordinates. We convert the bounding box to a normalized center point (range 0-1). This approach reduces the input context for LMMs while preserving spatial information.
3.2 RDAgent for semi-automated annotation
RDAgent (referring drone annotation framework with multi-agent system) is a semi-automated annotation framework that enhances traditional annotation pipelines by integrating LMMs with human annotators. Central to RDAgent are multiple LMM agents, specifically GPT-4o, each perform specialized tasks via distinct system prompts. As shown in Figure 2, RDAgent employs five structured steps.
Step 1: scene understanding. A GPT-4o-based captioning agent generates three diverse textual descriptions per image. These captions provide contextual perspectives (e.g., object relationships, spatial layouts) to anchor subsequent referring expression generation.
Figure 3: Object size distribution analysis. (a) Object size distribution in RefDrone dataset (small: $<32^{2}=1024$ pixels, normal: 1024 to 9216 pixels, large: $>96^{2}=9216$ pixels). (b) Object size histograms in RefDrone and gRefCOCO [22] datasets.
Step 2: color categorization. Color attributes serve as crucial discriminative features in referring expressions. Color attributes are extracted using a hybrid pipeline: (1) a WideResNet-101 [46] model trained on HSV color space-based labels, refined via manual validation, and (2) a dedicated LMM agent that verifies predictions to mitigate illumination- or occlusion-induced errors.
Step 3: expression generation. To address the complexity in referring expression generation, we reformulate the referring expression generation task as an object grouping problem. The target of the agent is to group semantically related objects and to provide appropriate reasons for each group. The reasons serve as the referring expressions. Furthermore, a dynamic feedback loop triggers color categorization (Step 2) when novel colors are detected, ensuring color attribute consistency.
Step 4: quality evaluation. An evaluator agent assesses the annotation semantic accuracy and referential uniqueness of object-expression pairs, categorizing annotations as:
  • “Yes.” The annotation is accurate and unique to the described objects.
  • “No.” The annotation is inaccurate or non-unique, with a detailed explanation of the discrepancy.
Annotations marked “Yes” proceed to Step 5. Annotations marked “No” activate specific feedback: expressions with semantic issues are returned to Step 3 (expression generation), while color-related inaccuracies are sent to Step 2 (color categorization) for refinement.
@@ 3.2 RDAgent for semi-automated annotation
  • Direct acceptance. Annotations satisfying all quality criteria are approved for the final dataset.
  • Refinement required. Annotations with minor errors are corrected through human editing.
  • Significant issues. Annotations with major errors or unclear descriptions are marked for complete reworking.
Annotations with significant issues enter a feedback loop, returning to Step 4 (quality evaluation). If an annotation repeatedly fails to meet standards, it becomes a no-target sample. These feedback refinement loops ensure that all annotations meet our quality standards and that no-target expressions still relate to the image content.
Each step utilizes LMMs through in-context learning with manually designed prompts and examples (see Supplementary Material). RDAgent requires only \$0.0539 pre expression in GPT4-o API cost and reduces human annotation effort by 88% compared to fully manual annotations. This cost-performance balance makes the framework scalable for large-scale datasets.
@@ 3.3 Dataset analysis
The RefDrone dataset contains 17,900 referring expressions across 8,536 images, comprising 63,679 object instances in 10 categories. We maintain the original train, validation, and test splits from VisDrone2019-DET [49]. The dataset features an average expression length of 9.0 words, with each expression referring to an average of 3.8 objects. Figure 1 gives several examples that relate to the three key features and challenges in RefDrone:
1) Multi-scale and small-scale target detection. Figure 3 presents an analysis of object scale distribution in RefDrone, with (a) showing the overall distribution and (b) offering a comparison with the gRefCOCO [22] dataset. In the RefDrone dataset, small-scale objects constitute 31% of all instances, while large-scale objects account for 14%. The detailed scale distribution in RefDrone demonstrates greater variance compared to gRefCOCO, highlighting the multi-scale challenges in RefDrone.
2) Multi-target and no-target samples. The RefDrone dataset includes 11,362 multi-target and 847 no-target expressions, with the number of referred targets ranging from 0 to 242. Figure 4 illustrates the target number distribution, revealing higher complexity in multi-target scenarios compared to gRefCOCO, where expressions typically refer to no more than two objects.
3) Complex environment with rich contextual expressions. The images present inherent complexity including diverse viewpoints, varying lighting conditions, and complex background environments. The referring expressions go beyond simple object attributes (e.g., color, category) and spatial relationships (e.g., ‘left’, ‘near’), incorporating rich object-object interactions (e.g., ‘the white trucks carrying livestock’) and object-environment interactions (e.g., ‘the white cars line up at the intersection’). Figure 5 presents word cloud visualizations of (a) complete expressions and (b) background terms, highlighting the diverse vocabulary used in descriptions.

### v2

@@ Abstract
Drones have become prevalent robotic platforms with significant potential in Embodied AI. A crucial capability for drone-based Embodied AI is Referring Expression Comprehension (REC), which enables locating objects with language expressions. Despite advances in REC for ground-level scenes, drones’ unique capability for broad observation introduces distinct challenges: multiple potential targets, small-scale objects, and complex environmental contexts. To address these challenges, we introduce RefDrone, an REC benchmark for drone scenes. RefDrone reveals three key challenges: 1) multi-target and no-target scenarios; 2) multi-scale and small-scale target detection; 3) complex environment with rich contextual reasoning. To efficiently construct this dataset, we develop RDAnnotator, a semi-automated annotation framework where specialized modules and human annotator collaborate through feedback loops. RDAnnotator ensures high-quality contextual expressions while reducing annotation costs. Furthermore, we propose Number GroundingDINO (NGDINO), a novel method to handle multi-target and no-target cases. NGDINO explicitly estimates the number of objects referred to in the expression and incorporates this numerical pattern into the detection process. Comprehensive experiments with state-of-the-art REC methods demonstrate that NGDINO achieves superior performance on RefDrone, as well as on the general-domain gRefCOCO and remote sensing RSVG benchmarks. The data and code are completely available at https://github.com/sunzc-sunny/refdrone.
@@ 1 Introduction
Drones/UAVs have become increasingly popular in our daily lives, serving both personal and professional purposes, such as entertainment, package delivery, traffic surveillance, and emergency rescue [1, 42, 45]. Their ability to move freely and to observe broadly make them important platforms for Embodied AI applications [12, 35, 23, 14, 31]. A crucial capability in Embodied AI is Referring Expression Comprehension (REC) [43, 48, 4, 49], which serves as a critical bridge between natural language understanding and visual perception. REC requires drones to localize specific objects in images based on natural language expressions. However, existing REC datasets primarily focus on ground-level perspectives, such as RefCOCO [54, 36] dataset. The application of REC in drone-based scenarios presents unique challenges.
@@
In this work, we introduce RefDrone, a challenging REC benchmark for drone scenes. RefDrone comprises 17,900 referring expressions annotated over 8,536 images, with 63,679 object instances. As illustrated in Figure 1, RefDrone poses three primary challenges that distinguish it from existing datasets: (1) multi-target and no-target samples, where expressions may refer to any number of objects, ranging from 0 to 242; (2) multi-scale and small-scale target detection, with 31% small objects and 14% large objects; and (3) complex environment with rich contextual reasoning, encompassing diverse viewpoints, lighting conditions, and intricate backgrounds. The referring expressions richly capture spatial relations, object attributes, and inter-object interactions. As summarized in Table 1, RefDrone offers greater diversity and complexity than previous REC benchmarks. We further evaluate 13 representative REC models using both fine-tuning and zero-shot settings. Notably, all models exhibit significantly poorer performance on RefDrone compared to standard REC datasets (e.g., GroundingDINO [34] yields 21.07 $Acc_{img.}$ on RefDrone vs. 91.4 $Acc_{img.}$ on RefCOCO), highlighting the inherent difficulty and unique challenges posed by our benchmark.
To efficiently construct RefDrone, we develop RDAnnotator, a semi-automated annotation pipeline for referring expression annotation in drone scenes. RDAnnotator leverages multiple specialized LMM-based modules, which collaborate within a feedback loop to generate and validate annotations. This tool reduces human involvement to quality control and minor adjustments, thereby significantly lowering annotation costs, while maintaining high-quality standards for complex expressions. RDAnnotator achieves a cost of just \$0.0539 per expression (GPT-4o API usage) and reduces human annotation time to under one minute per expression. This cost-efficiency makes RDAnnotator a scalable solution for large-scale dataset construction and can be readily adapted to other REC tasks.
@@
Furthermore, we propose Number GroundingDINO (NGDINO) to address the challenges of multi-target and no-target referring expressions. Our key insight is that explicitly utilizing the number information of referred objects enhances the handling of these scenarios. NGDINO includes three key components: (1) a number prediction head that estimates the count of target objects, (2) learnable number-queries that capture numerical patterns corresponding to varying object quantities, and (3) a number cross-attention module that fuses number queries with detection queries to inform object localization. Extensive experiments on both our RefDrone benchmark, as well as the general-domain gRefCOCO [28] and remote sensing RSVG [56] datasets, demonstrate that NGDINO substantially improves performance, particularly in the challenging multi-target and no-target cases.
@@
  • RefDrone Benchmark: We introduce RefDrone, a comprehensive benchmark for referring expression comprehension in drone scenes. RefDrone poses three key challenges: multi-target/no-target scenarios, multi-scale/small-object detection, and complex contextual reasoning. We provide thorough baseline evaluations with 13 representative REC models.
  • RDAnnotator Framework: We propose RDAnnotator, a semi-automated annotation framework that significantly reduces human effort while ensuring high-quality referring expressions. This framework is scalable for large-scale dataset construction and generalizable beyond drone imagery.
  • NGDINO Method: We develop NGDINO, a novel approach that explicitly models the number of referred objects to address multi-target and no-target cases. NGDINO achieves state-of-the-art performance on RefDrone and shows consistent improvements on gRefCOCO and RSVG benchmarks.
Table 1: Comparison of REC datasets relevant to RefDrone. Avg. objects: mean objects per expression. Avg. length: mean words per expression. No target: expressions without referred objects. Expression type: method for expression generation. Small target: percentage of small-scale objects.
RefCOCO/+/g [54, 36] gRefCOCO [28] D3 [51] RIS-CQ [19] RSVG [56] RefDrone (Ours) Image source COCO [26] COCO [26] COCO [26] VG [22]+COCO [26] DIOR [24] VisDrone [60] Avg. objects 1.0 1.4 1.3 3.6 2.2 3.8 Avg. length 3.6/3.5/8.4 4.9 6.3 13.2 7.5 9.0 No target Expression type Manual Manual Manual LLM (GPT3.5-t) Templated LMM (GPT4-o) Small target 0/0/0% 0.1% 6.3% - 17.2% 31.1%
@@ 2.1 Referring expression understanding datasets
Referring expression understanding identifies regions in visual content from natural language expressions. The main subtasks are Referring Expression Comprehension (REC), which outputs detection bounding boxes, and Referring Expression Segmentation (RES), which outputs segmentation masks. Datasets in this domain show increasing visual and linguistic complexity. Early datasets such as ReferIt [21] and RefCOCO [54] are pioneering but limited to simple expressions and single-target scenarios. Later datasets [32, 9, 28, 46, 19, 51] introduce more challenges. gRefCOCO [28] supports multi-target expressions. OV-VG [46] enables open-vocabulary queries. D3 [51] and RIS-CQ [19] offer more complex expressions. Domain-specific datasets [10, 8, 56, 37, 39] address specialized applications: RSVG [56] for remote sensing, RAVAR [37] for video action recognition, RIO [39] for affordance detection. However, drone scenes remain unexplored in REC, motivating our RefDrone benchmark. Table 1 compares RefDrone with other related REC benchmarks.
Recent dataset creation approaches leverage large language models (LLMs) to reduce annotation costs. Works such as LLaVA [30] and Ferret [53] demonstrate the effectiveness of LLMs in generating instruction tuning data. Similarly, RIS-CQ [19] and RIO [39] employ LLMs to generate referring expressions. However, these approaches treat LLMs as text generators and may lack proper visual grounding. In contrast, we employ a semi-automated annotation framework with iterative feedback mechanisms throughout the annotation pipeline to reduce annotation costs and improve quality.
@@ 2.2 Referring expression comprehension methods
REC methods can be broadly categorized into large multimodal models (LMMs) and specialist models. Large Multimodal Models [30, 53, 5, 7, 3, 38, 47, 27, 57, 50] have recently been applied to REC tasks as part of evaluating their broader visual-language understanding capabilities. These models leverage extensive referring instruction tuning data to achieve competitive performance without task-specific architectural designs. Despite their broad capabilities, LMMs consistently struggle with small object detection, primarily due to their inherent input resolution constraints. This constraint forces image downsampling, resulting in the loss of fine-grained visual details.
Specialist models include two-stage and one-stage methods. Two-stage methods [18, 29, 16, 44, 15] typically approach REC as a ranking task: first generating region proposals through a detector, then ranking them based on their alignment with the language input. However, they are often criticized for slow inference speeds. In contrast, one-stage methods [20, 25, 33, 13, 52, 34] directly predict target regions guided by language input. These approaches leverage transformer to enable cross-modal interactions between visual and textual features. Among these methods, GroundingDINO (GDINO) [34] has gained widespread attention for its impressive results in REC tasks. Our work extends GDINO by incorporating explicit number modeling to handle multi-target and no-target scenarios, which are critical challenges in drone-based REC tasks.
@@ 3 RefDrone benchmark
Figure 2: The overview of the RefDrone annotation process with RDAnnotator. Multiple specialized LMM-based modules collaborate both with each other and human annotators through iterative feedback loops to generate high-quality annotations.
@@ 3.1 Data source
The RefDrone benchmark is built upon VisDrone2019-DET [60], a large-scale high-quality drone-captured object detection dataset. Source images are collected across multiple scenarios, illumination conditions, and flying altitudes. To ensure meaningful visual content, we filter out images with fewer than three objects and objects with bounding box areas below 64 pixels. VisDrone2019-DET provides object categories and bounding box coordinates, which we convert to normalized center points (range 0-1). This approach reduces the input context for LMMs while preserving spatial information.
3.2 RDAnnotator for semi-automated annotation
RDAnnotator (referring drone annotator) is a semi-automated annotation framework that minimizes human effort while ensuring high quality. It integrates specialized modules powered by LMMs with human validation. As shown in Figure 2, RDAnnotator employs five structured steps.
Step 1: scene understanding. A GPT-4o-based scene parsing module generates three diverse textual descriptions per image. These captions establish spatial layouts, object relationships, and contextual foundations to guide subsequent referring expression generation.
Step 2: color categorization. Color attributes are crucial discriminative features in referring expressions. We implement a hybrid color extraction pipeline, which combines low-level CNN-based classification (WideResNet-101 classifier [55]) with high-level LMM semantic reasoning, ensuring robust color attribution in challenging aerial conditions.
Step 3: expression generation. To address the complexity of referring expression generation, we reformulate the referring expression generation task as an object grouping problem. The target of the module is to group semantically related objects and provide appropriate reasons for each group. The reasons serve as the referring expressions. Furthermore, a dynamic feedback loop triggers Step 2 (color categorization) when unseen colors are detected, ensuring color attribute consistency.
Step 4: quality evaluation. A validation module assesses the semantic accuracy and uniqueness of each object-expression pair, classifying each annotation as either correct (accurately and uniquely identifying the described objects) or incorrect (inaccurate or ambiguous, with an explanation of the identified issue). Annotations deemed correct proceed to Step 5 (human verification), while incorrect annotations trigger targeted feedback: those with semantic errors are returned to Step 3 (expression generation), and color-related inaccuracies are redirected to Step 2 (color categorization).
Figure 4: Object size distribution histograms in RefDrone and gRefCOCO [28] datasets.
@@ 3.2 RDAnnotator for semi-automated annotation
  • Direct acceptance. Annotations satisfying all criteria are approved for the final dataset.
  • Refinement required. Minor errors are corrected through human editing.
  • Significant issues. Poorly grounded or inconsistent annotations trigger full regeneration.
Annotations with significant issues re-enter the pipeline at Step 4 (quality evaluation). If an annotation repeatedly fails, it is classified as a no-target sample. These feedback loops ensure all annotations meet quality standards and that no-target expressions are contextually relevant to the image.
Each step utilizes LMMs through in-context learning with carefully designed task-specific prompts and examples (see Appendix A.8). The final annotation quality demonstrates 42% directly accepted samples, 47% requiring minor refinement, and 11% demanding full re-annotation. RDAnnotator reduces human annotation effort by 85%, decreasing annotation time from 7 to 1 minute per expression, while incurring only \$0.0539 per expression in GPT-4o API costs. This cost-performance balance makes the framework scalable for large-scale dataset creation while maintaining annotation quality.
@@ 3.3 Dataset analysis
The RefDrone dataset contains 17,900 referring expressions across 8,536 images, comprising 63,679 object instances in 10 categories. We retain the original train, validation, and test splits from VisDrone2019-DET [60]. The average expression length is 9.0 words, and each expression refers to an average of 3.8 objects. Figure 1 highlights the three key challenges in RefDrone:
1) Multi-target and no-target samples. Unlike conventional REC datasets [54, 36] that primarily focus on single-object references, RefDrone includes 11,362 multi-target and 847 no-target expressions. The number of referred targets per expression ranges from 0 to 242. Figure 4 illustrates the target number distribution, revealing higher complexity in multi-target scenarios compared to gRefCOCO, where the majority of expressions contain references to only one or two objects.
2) Multi-scale and small-scale target detection. Figure 4 presents the challenges of scale distribution: small-scale objects ($<32^{2}$ pixels) make up 31% of instances, normal-scale ($32^{2}$-$96^{2}$ pixels) 55%, and large-scale ($>96^{2}$ pixels) 14%. Compared to gRefCOCO [28], RefDrone has a notably higher variance in object scales, emphasizing the dataset’s multi-scale challenges.
3) Complex environment with rich contextual reasoning. The images present inherent complexity, including diverse viewpoints, varying lighting conditions, and complex background environments. The referring expressions go beyond simple object attributes (e.g., color, category) and spatial relationships (e.g., ‘left’, ‘near’), incorporating rich object-object interactions (e.g., ‘the white trucks carrying livestock’) and object-environment interactions (e.g., ‘the white cars line up at the intersection’). Figure 5 visualizes this linguistic richness through word clouds of (a) complete expressions and (b) background terms, highlighting the diverse vocabulary used in descriptions.